OpenStudy (anonymous):

simplify the expression, 5x^4-10x^3-15x^2/7x^2

OpenStudy (anonymous):

(5x^4-10x^3-15x^2)/(7x^2) ==> 5/7(x+1)(x-3) ; 5/7(x-3)(x+1) ; 5/7(x^2-2x-3)

OpenStudy (anonymous):

did you factor the numerator?

OpenStudy (anonymous):

5x^4-10x^3-15x^2 => 5x^2(x^2-2x-3) The x^2 will cancel out with the x^2 in 5x^2

OpenStudy (anonymous):

Then you can play around with (x^2-2x-3) in anyway you like

OpenStudy (anonymous):

It's not clear to me please.

OpenStudy (anonymous):

\[\frac{5}{7} (x-3) (x+1) \]

OpenStudy (anonymous):

why the 5/7@robtobey

OpenStudy (anonymous):

\[\frac{5 x^4-10 x^3-15 x^2}{7 x^2}=\frac{5 x^2}{7}-\frac{10 x}{7}-\frac{15}{7}=\frac{5}{7} (x-3) (x+1) \]

OpenStudy (anonymous):

ok...which did you factor to get (x-3)(x+1)

OpenStudy (anonymous):

\[\frac{5}{7} \left(x^2-2 x-3\right) \]

OpenStudy (anonymous):

@robtobey ...please let me understand how you got (x^2-2x-3)

OpenStudy (whpalmer4):

\[\frac{5x^2}{7}-\frac{10x}{7} -\frac{15}{7} = \frac{5}{7}(x^2-2x-3)\]

OpenStudy (anonymous):

I don't understand where x^2-2x-3 came from...pls can somebody explian please

OpenStudy (anonymous):

\[\frac{1}{7} \left(5 x^2-10 x-15\right) \]Factor out 5 from the numerator of the above.

OpenStudy (anonymous):

\[5 x^2-10 x-15=5 \left(x^2-2 x-3\right) \]

OpenStudy (anonymous):

x^2-2x-3

OpenStudy (anonymous):

ok..now i know how...thanks guys

OpenStudy (anonymous):

You're welcome.